OpenStudy (anonymous):

Is my PV=nRT question correct? 'mixture of hydrogen and oxygen at 101000 Pa and 0 degrees C occupies 200cm^3, What are total number of moles?

OpenStudy (anonymous):

I did n = PV/RT , 101000 * (200cm^3 goes to 0.0002m^3) = 20.2 Then 8.314 * 273.15 = 2270.961 So 20.2 / 2270.961 = 0.008894mol , should this be in KJMol and * by 1000?

OpenStudy (unklerhaukus):

the units for the gas constant are joules per mole per kelvin \(R=8.314 [\text J\, \text{mol}^{-1}\,\text K^{-1}]\)

OpenStudy (unklerhaukus):

the unit for the pressure was pascals? \([\text{Pa}]=[\text N/\text m^2]\)

OpenStudy (unklerhaukus):

a newton equals a joule / meter \([\text N]=\,[\text J/\text m]\)

OpenStudy (anonymous):

But is the answer correct however?

OpenStudy (chmvijay):

what ?

OpenStudy (anonymous):

0.008894mol ?

OpenStudy (chmvijay):

ur right :)

OpenStudy (anonymous):

Thanks, and the untis for the final answer? The conversion between cm^3 and m^3 ?

OpenStudy (chmvijay):